Cos x - 2sin^2x=1 find all real numbers that satisfy the equation

Answer by stanbon(75887)

Cos x - 2sin^2x=1 find all real numbers that satisfy the equation
Cos x - 2sin^2x = 1
cos(x) - 2(1-cos^2(x)) = 1
-----
cos(x) - 2 + 2cos^2(x) = 1
-----
Rearrange:
2cos^2(x) + cos(x) - 3 = 0
====
Factor::
(2cos(x) +3)((cos(x)-1) = 0
============
cos(x) = -3/2 or cos(x) = 1
----
Note:: cos(x)= -3/2 is extraneous.
------
x = 0
================
Cheers,
Stan H.
================